# Choi-Defined Resource Theories

### v1

@@ Abstract
In this work, we identified a key property of the resource theories of magic, imaginarity, SEP, and NPT, which is that a channel is considered free if and only if its suitably renormalized Choi matrix is a free state. We presented a new way to construct a resource theory from the set of free states based on Choi matrices and we derived which conditions on the free states are necessary and sufficient for constructing a well-defined resource theory. Our interest is driven by the significant advantages that such a property guarantees. These advantages stem from the fact that every question in Choi-defined resource theories can be naturally framed as a question involving free states alone, which often have a well-understood structure. Moreover, we showed that in any CDRT free operations are all CRNG operations. This provides a new construction for CRNG operations when the set of free states satisfy the conditions of Theorem 3. A natural future direction is to explore Choi-defined dynamical resource theories [55, 56, 57], i.e., resource theories for channels, where a superchannel [58, 59] is free if and only its Choi matrix is the Choi matrix of a free channel. It is noteworthy that such property has already been considered in the resource theories for dynamical entanglement [46] and magic [60], where it led to interesting results. The characterization of Choi-defined dynamical resource theories is a particularly challenging problem because the techniques employed in this work cannot be easily translated for superchannels. Indeed, there are some subtleties in the definition of the Choi matrix of a superchannels. For example, in Ref. [58], several different Choi matrices are constructed from the same superchannel, which differ in the order systems are considered. As in the case studied in this Letter, an ordering of systems is mandatory if one wants to construct resource theories from Choi matrices, as inputs and outputs play different roles especially if we want such construction to be generalizable, without ambiguity, to even higher-order maps [61, 62].
@@ Acknowledgements.
The authors are grateful to Matt Wilson for the several in-depth discussions on the subject of this letter and to Gilad Gour for the interesting conversations about this research. E. Z. acknowledges support from the Alberta Graduate Excellence Scholarship (AGES) and from the Eyes High International Doctoral Scholarship. C. M. S. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grant “The power of quantum resources” RGPIN-2022-03025 and the Discovery Launch Supplement DGECR-2022-00119.

### v2

@@ Abstract
In this work, we identified a key property of the resource theories of magic, imaginarity, SEP, NPT, and non-negativity of quantum amplitudes, which is that a channel is considered free if and only if its suitably renormalized Choi matrix is a free state. We presented a new way to construct a resource theory from the set of free states based on the Choi isomorphism and we derived which conditions on the free states are necessary and sufficient for constructing a mathematically consistent Choi-defined resource theory. Our interest is driven by the significant advantages that such a property guarantees. These advantages stem from the fact that every question in Choi-defined resource theories can be naturally framed as a question involving free states alone, which often have a well-understood structure. Indeed, we introduce several resource quantifiers which can all be computed with CLPs if the set of free states is convex and closed. Moreover, we showed that in any CDRT free operations are all CRNG operations. This provides a new constructive way to define CRNG operations when the set of free states satisfy the conditions of Theorem 3.
A natural future direction is to explore Choi-defined dynamical resource theories [61, 63, 64], i.e., resource theories for channels, where a superchannel [65, 66] is free if and only its Choi matrix is the Choi matrix of a free channel. It is noteworthy that such property has already been considered in the resource theories for dynamical entanglement [46] and magic [67], where it led to interesting results. The characterization of Choi-defined dynamical resource theories is a particularly challenging problem because the techniques employed in this work cannot be easily translated for superchannels. Indeed, there are some subtleties in the definition of the Choi matrix of a superchannels. For example, in Ref. [65], several different Choi matrices are constructed from the same superchannel, which differ in the order systems are considered. As in the case studied in this Letter, an ordering of systems is mandatory if one wants to construct resource theories from Choi matrices, as inputs and outputs play different roles especially if we want such construction to be generalizable, without ambiguity, to even higher-order maps [68, 69].
@@ Acknowledgements.
The authors are grateful to Matt Wilson for the several in-depth discussions on the subject of this letter and to Gilad Gour and Nathaniel Johnston for the interesting conversations about this research. E. Z. acknowledges support from the Alberta Graduate Excellence Scholarship (AGES) and from the Eyes High International Doctoral Scholarship. C. M. S. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grant “The power of quantum resources” RGPIN-2022-03025 and the Discovery Launch Supplement DGECR-2022-00119.